OpenStudy (anonymous):

Determine if the sequence is arithimetic if it is find the common difference . 1. 13 , 7, 1 , -5 A. d=-4 B. d=-2 C.d=2 D. d=-6

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen please help

OpenStudy (cwrw238):

subtract the second term from the first, the third from the second and the fourth from the third. What values do you get?

OpenStudy (anonymous):

Okay i will

OpenStudy (anonymous):

5

OpenStudy (cwrw238):

oops sorry i meant second - first, etc

OpenStudy (cwrw238):

7 - 13= 1 - 7 = -5 - 1 =

OpenStudy (anonymous):

-6 sir

OpenStudy (cwrw238):

right -6 for each subtraction thats your common difference it is an arithmetic series

OpenStudy (anonymous):

Okay thank you can you please help me with others

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

okay ill tag you